xx IDConverterILGRMHD: A Module for Converting HydroBase Initial Data into Variables IllinoisGRMHD Can Read

IDxConverterxILGRMHD: A Module for Converting HydroBase

Initial Data into Variables IllinoisGRMHD Can Read

Original author: Zachariah B. Etienne.

Date: 2015-10-12 12:00:00 -0600 (Mon, 12 Oct 2015)

Abstract

IllinoisGRMHD and HydroBase variables are incompatible; The former uses 3-velocity defined as , and the latter uses the Valencia formalism definition of .

Define the Valencia formalism’s definition of to be ””, and IllinoisGRMHD’s definition ”” Then

which comes from Eq 11 in the IllinoisGRMHD code announcement paper:

http://arxiv.org/pdf/1501.07276.pdf.

Similarly,

In addition, IllinoisGRMHD needs the A-fields to be defined on staggered grids, and HydroBase does not yet support this option. The staggerings are defined in Table 1 of the IllinoisGRMHD code announcement paper:

http://arxiv.org/pdf/1501.07276.pdf (page 15).

The long-term goal should be to adjust HydroBase and IllinoisGRMHD so that this thorn is no longer necessary.

1 Parameters

11gammaxinitial10Scope: privatex01  REAL

11Description: Single Gamma-law EOS: Gamma

11 Range 10 x01 Default: 1.3
110:*  10Positive or zero

11

11kxinitial  10Scope: privatex01  REAL

11Description: Single Gamma-law EOS: K

11 Range 10 x01 Default: 1.0
110:*  10Positive or zero

11

11purexhydroxrun 10Scope: privatex01  BOOLEAN

11Description: Set the vector potential and corresponding EM gauge quantity to zero

11            10  x01  Default: no

11

11randomxpert 10Scope: privatex01  REAL

11Description: Random perturbation atop data

11 Range 10 x01 Default: (none)
11*:*  10Anything goes.

11

11randomxseed 10Scope: privatex01  INT

11Description: Random seed for random, generally roundoff-level perturbation on initial data. Seeds srand(), and rand() is used for the RNG.

11 Range 10  x01  Default: (none)
110:99999999 10Anything unsigned goes.

11

2 Interfaces

General Implements: idxconverterxilgrmhd 3 Schedule This section lists all the variables which are assigned storage by thorn WVUThorns/IDxconverterxILGRMHD. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage NONE